Bob deposited $409274 into an account paying 2% interest compounded annually with the goal of letting the money accrue interest until the end of year 34. Unfortunately, Bob had a small health crisis at the end of year 8 and had to use the money in the account for several years to help with expenses. Bob made an annual withdrawal of $30128 at the end of each year for 4 years. Bob's health improved and he was able to let the remainder of the money accrue interest as he had originally planned. After the emergency withdrawals, how much was left in the account at the end of the 34 years?
Answer #1

Present value of withdrawal = 30128*(P/A,2%,4) * (P/F,2%,7)

= 30128 * 3.807729 * 0.87056

= 99870.01

NPW = 409274 - 99870.01 = 309403.99

Future worth = 309403.99 * (F/P, 2%,34) = 309403.99 * 1.960676 = 606640.98 = 606641 (rounding off)